1. points A, B, and C are collinear and A is between B and C.
AB= 4x -3 BC =7x + 5, and AC = 5x -16
Find each value.
a. BC
b. AB
c. AC
Please e3xplain how to solve this problem

Found 2 solutions by solver91311, rameens_99:
Answer by solver91311(24713)



If A, B, and C are collinear and A is between B and C, then



So:



Solve for , then substitute back into each of the expressions for BC, AB, and AC to calculate the three required measures.

John

My calculator said it, I believe it, that settles it


Answer by rameens_99(10)

BA= 4x-3
AC= 5x-16
BC= 7x+5


BA+AC= BC


4x-3+5x-16=7x+5
9x-3-16=7x+5
9x-19=7x+5
9x=7x+24
2x=24
____________
x=12




Now plug it in each equation in place of x!!!
BA= 4x-3
4(12)-3
48-3
__________
BA=45


AC= 5x-16
5(12)-16
60-16
______________
AC=44



BC= 7x+5
7(12)+5
84+5
____________
BC=89





You're Welcome (-: